# Designing Robots to Help Women*

### v1

@@ V DISCUSSION
In the current paper, three main categories of challenges that disproportionately affect women were identified: criminal, medical, and day-to-day. From these challenges were born ideas for opportunities, of how designers could use robots to better women’s lives. Five scenarios were visualized via sketches. Furthermore, the first sketch was implemented as a prototype, of a drone system that uses a thermal/RGB camera to detect hidden cameras. A summary of this work is also available via an online video. [^23] https://youtu.be/7slpfGD1sEU
@@
Regarding the simplified prototype intended to detect hidden cameras, the result of 80% accuracy (IoU 0.401) seemed reasonable, given the challenging mock-up and the simplified approach. As a result of some simplified experimentation, we observed that lowering the confidence parameter allows YOLO to detect more bounding boxes, such that all five objects of interest are detected; a demerit is that reducing the confidence threshold results in many more objects detected, which can result in overlapping bounding boxes and increased complexity. Also, although the IoU score is imperfect, we believe this should not be a deal-breaker since a robot or person is not limited to checking only the exact inside of each bounding box but can also check the vicinity, as long as they have a rough estimate of where the camera might be.
@@
Also, YOLO can be used not only to obtain bounding boxes, but also to recognize objects. This could be useful in various ways. For example, some natural heat sources like ovens or people could be removed from consideration. For others, such as lamps or computers, a model of where such devices normally become warm could be used to detect anomalies caused by hidden cameras. As well, a label could be directly provided to the human in regard to which objects are expected to contain hidden cameras. However, in our simplified exploration, object recognition appeared to have been challenging, possibly due to illumination, the cluttered mock-up environment, or camera limitations. As can be seen in Fig. 4, four objects were recognized correctly, two teddy bears, a clock, and a book, whereas seven objects were recognized incorrectly, as a sink, cup, bottle, donut, cellphone, knife, and book. In such a case, a robot might be able to avoid having to recognize objects by instead showing people where cameras are expected to be located on a screen (e.g., either a screen on the robot, or by sending a picture to a person’s smartphone).
@@
Along the side, one thing we observed during preparation was that hooking up all five cameras to an old desktop resulted in crashes, possibly due to overwhelming the bandwidth on the same USB bus, so for subsequent attempts several computers were used. Another observation was that our drone, like other typical drones used by the community such as Crazyflie, would not able to detect objects on the ceiling due to its camera placements, as it only has two cameras, that look forward and below respectively.
V-A Limitations and Future Work
The current work is limited by various factors, including a bias toward focusing on Sweden, given our location. As well, our ideas come only from us, whereas future studies should gather ideas from a larger pool of female participants, in different countries. Our ideas are also speculative, and thus by intention largely ignore some current realities: e.g., that robots are costly, difficult to maintain and fix, and limited in battery life and capabilities. For the prototype, results are limited by the controlled lab environment (e.g., detecting of expensive, upper range spy cameras from blurry images of a moving drone might require more complex methods).
@@ V-A Limitations and Future Work
Furthermore, new technologies create not only opportunities, but also potentially new crimes. Robot designers should consider ahead of time how robots could be misused, and factor in such questions. For example, how to ensure that times or images recorded by a robot are accurate. Another question is if robots could inadvertently harm the women that they seek to help. For example, wind from a drone’s propellers could disturb a crime scene, stirring up dust, erasing evidence or introducing contamination. Or, robots could make it harder for women to escape or hide (e.g. if a costly or large robot crashes, blocks an escape route, or betrays a woman’s location). As well, the concepts in the developed sketches could also be potentially misused to harm women. For example, the camera detecting prototype could be used by criminals to ensure that their cameras are well-hidden and will not be easily detected. Or, a robot that guards a woman’s personal space could be hacked to slip a sedative into her drink. Furthermore, (although it is estimated that false accusations, e.g., of rape, tend to be not so common) criminals could also potentially misuse robots to get others in trouble (e.g., getting an accomplice wearing a mask to trick the robot into thinking an innocent person committed a crime). Thus, one important area of future work will involve refining such scenarios, such that for example robots could be used to facilitate legal processes by delivering new or better forms of evidence in courts.
@@
Another question regards who will control women’s ability to protect themselves with robots. Although governments introduce various rules and regulations, they might not "set the pace", since criminals can ignore rules, and people often use apps, which have their own terms. While users cannot be aware of all sets of terms for the apps they use, even if rules of technology providers might be in conflict, it seems like women themselves could set up the apps and robots they need for protection, e.g., detecting rapes in real-time with a drone. And, if such usage becomes widespread, it could motivate modifying existing restrictions. Therefore, we believe that prototyping solutions in this area could be a useful showcase for what AI can do to help people.
@@
In summary, we believe that further exploration of the challenges and opportunities that arise in designing robots to help women could help us to realize a safer, more inclusive future for women, and thereby for all of society.

### v2

@@ V DISCUSSION
The current paper has sought to highlight an important but little-addressed topic: millions of women face problems that could be mitigated via robotics. Three categories of challenges that disproportionately affect women were identified, in relation to crimes, health, and daily activities–comprising specific challenges such as voyeurism, DV, rape, difficult pregnancies, physical inequalities and unpaid care work. From these challenges were born ideas of how designers could use robots to better women’s lives by interfering with crimes (deterring and bringing to justice), democratizing health care (detecting and intervening), and equalizing opportunities (physically and time-wise). Five ideas were visualized via concrete sketched examples, illustrating how flying, wheeled, humanoid, and exoskeletal robot designs could aim to help young women seeking privacy and safety, pregnant women, mothers, and working women. Furthermore, the first sketch was implemented as a prototype, of a drone system that uses a thermal/RGB camera to detect hidden cameras. A summary of this work is also available via an online video. [^28] https://youtu.be/7slpfGD1sEU
V-A Reflections on the Prototype
@@
Regarding the simplified prototype intended to detect hidden cameras, the result of 80% accuracy (IoU 0.401) seemed reasonable, given the challenging mock-up environment and simplified detection approach. We also observed that lowering the confidence parameter allows YOLO to detect more bounding boxes, such that all five objects of interest are detected; a demerit is that this results in many more objects detected, and thereby a higher risk of overlapping bounding boxes and increased complexity. Furthermore, although the IoU score is imperfect, we believe this should not be a deal-breaker since a robot or person is not limited to checking only the exact inside of each bounding box but can also check the vicinity, if a rough estimate is available of where a camera might be.
@@
Moreover, YOLO’s object recognition capability could also be used: For example, some alternative heat sources like ovens or people could be removed from consideration; other devices like lamps or computers could be checked for anomalous heat patterns using a "normal" model. As well, a label could be directly provided to a human in regard to which objects are expected to contain hidden cameras. However, in our simplified exploration, object recognition appeared to have been challenging, possibly due to illumination, the cluttered mock-up environment, or camera limitations. As can be seen in Fig. 4, four objects were recognized correctly–two teddy bears, a clock, and a book–whereas seven objects were recognized incorrectly–as a sink, cup, bottle, donut, cellphone, knife, and book. This problem could be avoided by instead showing people where cameras might be located (e.g., either using a screen on the robot, or by sending a picture to a person’s smartphone).
@@
Along the side, one thing we observed during preparation was that hooking up all five cameras to an old desktop resulted in crashes, possibly due to overwhelming the bandwidth on the same USB bus, so for subsequent attempts several computers were used. Another observation was that our drone, like other typical drones used by the community, would not able to detect objects on the ceiling due to its camera placements, as it only has two cameras, that look forward and downward respectively.
V-B Limitations and Future Work
The current work is limited by its exploratory nature: involving a small group of experts focused on Sweden, and ignoring current practical limitations of robots. Future studies can gather ideas from a larger pool of female participants with different backgrounds, or automatically identify challenges from the literature using AI methods, and consider factors such as cost, maintenance, battery life, and capabilities. Moreover, sketches are examples and not "only alternatives": for example, a wheeled robot with long battery life that can hold heavy, powerful sensors could be used in place of a drone if visibility of high-up places is not required. For the prototype, results are also limited by the controlled lab environment (e.g., detecting expensive, upper-range spy cameras in blurry images from a moving drone might require more complex methods). Furthermore, the current paper, which falls in the area of HRI and robot design, mostly does not offer technical details of how solutions should be implemented–instead focusing on what we felt was the first fundamental problem, of obtaining a "lay of the land". (We note too that the paper should not be interpreted as claiming that women’s needs should be prioritized over the needs of others; rather the aim is to expose some new challenges whose solutions could benefit all of society.)
@@ V-B Limitations and Future Work
In addition to further developing the drone prototype and prototyping the other sketches, future work will explore potential threats and regulation: Given that new technologies create not only opportunities, but also potentially new problems, robot designs should factor in ahead of time how robots could fail or be misused; e.g., how to ensure that times or images recorded by a robot are accurate. Another question is if robots could inadvertently harm the women they seek to help. For example, wind from a drone’s propellers could disturb a crime scene, stirring up dust, erasing evidence or introducing contamination. Or, robots could make it harder for women to escape or hide (e.g., if a costly or large robot crashes, blocks an escape route, or betrays a woman’s location). As well, the concepts in the developed sketches could also be potentially misused to harm women. For example, a camera-detecting robot could be used by criminals to test that their cameras are well-hidden and will not be easily detected. Or, a robot that guards a woman’s personal space could be hacked to slip a sedative into her drink. Furthermore, criminals could also potentially misuse robots to get others in trouble (e.g., using disguises to trick a robot into thinking an innocent person committed a crime). Thus, one important area of future work will involve refining such designs, such that, e.g., robots could facilitate legal or healthcare processes by delivering new or better forms of evidence in courts or hospitals.
@@
Another central question regards who will control women’s ability to protect themselves with robots. Although governments introduce various rules and regulations, these might not "set the pace", since criminals can ignore rules, and people often use apps, which have their own terms decided by technology providers. While users cannot be aware of all sets of terms for the apps they use, even if terms might be in conflict, it seems like women themselves could set up the apps and robots they need for protection, e.g., detecting rapes in real-time with a drone. And, if such usage becomes widespread, it could motivate modifying existing restrictions. Therefore, we believe that prototyping solutions in this area could be a useful showcase for what AI can do to help people.
@@
In summary, we believe that further exploration of how technologies can be used to help various groups, including women, will contribute to realizing a safer, healthier, more inclusive, and better future for all of society.